# Software paper for submission to the Journal of Open Research Software

### v1

@@ Abstract
PDSim is an R package that enables users to simulate commodity futures prices using the polynomial diffusion model introduced in Filipović and Larsson [7] through both a Shiny web application and R scripts. It also provides state variables and contract estimations via the Extended Kalman Filter (EKF) or Unscented Kalman Filter (UKF). With its user-friendly interface, PDSim makes the features of simulations and estimations accessible to all users. To date, it is the only package specifically designed for the simulation and estimation of the polynomial diffusion model. Additionally, the package integrates the Schwartz and Smith two-factor model [17] as an alternative approach. PDSim offers versatile deployment options, including running locally, via the Shiny server, or through Docker.
@@
Polynomial diffusion model; Schwartz-Smith model; commodity futures; two-factor modelling
@@ Introduction
Stochastic models are crucial in the analysis of commodity futures, serving important roles in various financial areas such as price forecasting, risk management, and portfolio optimisation. Typically, the underlying spot price, denoted as $S_{t}$, is modelled as a function of certain factors. In 1990, the Ornstein-Uhlenbeck (OU) process was introduced to model oil futures in a two-factor setup, representing spot price and convenience yield [8]. Building on this foundation, Schwartz and Smith [17] modelled the logarithm of the underlying spot price of crude oil futures as the sum of two hidden factors. These factors, assumed to follow the OU process, capture short-term fluctuations and the long-term equilibrium price level, respectively. This latent factor model and its extensions have been widely used in stochastic modelling in various fields, including biology [6], electricity forwards [13], agricultural commodity futures [18], and crude oil futures [2, 5, 16].
Despite its widespread use, this framework has certain limitations. First, a closed-form expression is required for futures prices. To achieve this, it is common to assume that all involved factors follow a Gaussian distribution and that the logarithm of the spot price is a linear function of these factors. This assumption ensures a log-normal distribution of the spot price. Second, within this framework, the spot price is typically modelled in logarithmic form, which generally works well. However, an unprecedented event on April 20, 2020, challenged this approach: the front-month May 2020 WTI crude oil futures settled at an extraordinary value of -\$37.63 per barrel on the New York Mercantile Exchange. This rare phenomenon significantly undermines the validity of the entire framework.
@@
The polynomial diffusion model was introduced to overcome these limitations. In this framework, the spot price is represented as a polynomial of any order in terms of the factors. In particular, under certain conditions, it can be proven that the conditional expectation of the spot price, which is equivalent to the futures price under the assumption of an arbitrage-free market and non-stochastic interest rate, is also a polynomial in terms of factors. The mathematical foundations of the polynomial diffusion model were introduced in [7], with applications of this framework seen in the modelling of electricity forwards, where the spot price is represented by a quadratic form of two factors [14].
I. Schwartz and Smith two-factor model
@@
where $\chi_{t}$ represents the short-term fluctuation and $\xi_{t}$ is the long-term equilibrium price level. Additionally, we assume both $\chi_{t}$ and $\xi_{t}$ follow a risk-neutral Ornstein-Uhlenbeck process,
@@
In the original Schwartz-Smith model [17], the long-term factor $\xi_{t}$ is geometric Brownian motion (gBm), corresponding to setting $\gamma=0$. Since 2000, many studies such as [2, 4, 15, 16] and others considered the extended version of the Schwartz-Smith model by introducing $\gamma\geq 0$. For simplicity, further, we continue to refer to this model with $\gamma\geq 0$ as the Schwartz-Smith model. Setting $\lambda_{\chi}=\lambda_{\xi}=0$ in Equation 2 and Equation 3 gives the real processes of the factors. The risk-neutral processes are used to calculate futures prices, and the real processes are for modelling state variables in real time.
@@
where $x_{t}=\left[\begin{matrix}\chi_{t}\\ \xi_{t}\end{matrix}\right],c=\left[\begin{matrix}0\\ \frac{\mu_{\xi}}{\gamma}\left(1-e^{-\gamma\Delta t}\right)\end{matrix}\right],E=\left[\begin{matrix}e^{-\kappa\Delta t}&0\\ 0&e^{-\gamma\Delta t}\end{matrix}\right],y_{t}=\left(\log{(F_{t,T_{1}})},\dots,\log{(F_{t,T_{m}})}\right)^{\top}$, $d_{t}=\left(A(T_{1}-t),\dots,A(T_{m}-t)\right)^{\top},F_{t}=\left[\begin{matrix}e^{-\kappa(T_{1}-t)},\dots,e^{-\kappa(T_{m}-t)}\\ e^{-\gamma(T_{1}-t)},\dots,e^{-\gamma(T_{m}-t)}\end{matrix}\right]^{\top}$ and $m$ is the number of futures contracts. The function $A(\cdot)$ is given by
@@
where $\tau$ is the time to maturity. $w_{t}$ and $v_{t}$ are multivariate Gaussian noises with mean 0 and covariance matrix $\Sigma_{w}$ and $\Sigma_{v}$ respectively, where
@@ II. Polynomial diffusion model
Under the polynomial diffusion framework, the spot price $S_{t}$ is expressed as a polynomial function of the hidden state vector $x_{t}$ (with components $\chi_{t}$ and $\xi_{t}$):
@@
In this context, we assume that the polynomial $p_{n}(x_{t})$ has a degree of 2 (or 1 if $\alpha_{4}=\alpha_{5}=\alpha_{6}=0$). However, it is worth noting that all the theorems presented here are applicable even for polynomials of higher degrees. Additionally, similar to the Schwartz and Smith model, we assume that $\chi_{t}$ and $\xi_{t}$ follow an Ornstein–Uhlenbeck process
@@
for $x\in\mathbb{R}^{d}$ and any $C^{2}$ function $f$. Let $N$ be the dimension of $Pol_{n}$, and $H:\mathbb{R}^{d}\to\mathbb{R}^{N}$ be a function whose components form a basis of $Pol_{n}$. Then for any $p\in Pol_{n}$, there exists a unique vector $\vec{p}\in\mathbb{R}^{N}$ such that
@@
Obviously, the hidden state vector $x_{t}$ satisfies the SDE with
@@
The Kalman Filter (KF) [9] is a commonly used filtering method in estimating hidden state variables. However, KF can only deal with the linear Gaussian state model. To capture the non-linear dynamics in the PD model, we use Extended Kalman Filter (EKF) [11] and Unscented Kalman Filter (UKF) [12, 19]. Suppose we have a non-linear state-space model
@@ IV. Comparison with existing libraries
The graphical user interface (GUI) provides an easy way for anyone to use the PDSim package, even without programming knowledge. Simply enter the necessary parameters, and PDSim will simulate the data and generate well-designed interactive visualizations. Currently, PDSim supports data simulation from two models: the Schwartz and Smith two-factor model [17] and the polynomial diffusion model [7]. In this section, we will guide you through using the GUI to simulate data.
@@ I. Implementation of the Schwartz-Smith model
First, we establish some global configurations, such as defining the number of observations (trading days) and contracts. Additionally, we select the model from which the simulated data will be generated.
Figure 4: Specify model parameters (Schwartz and Smith model).
For the Schwartz-Smith model, we assume that the logarithm of the spot price $S_{t}$ is the sum of two hidden factors:
@@
Here, $\kappa,\gamma\in\mathbb{R}^{+}$ are the speed of mean-reversion parameters, controlling how quickly these two latent factors converge to their mean levels. Most experiments suggest that $\kappa$ and $\gamma$ fall within the range $(0,3]$. To avoid issues with parameter identification, we recommend setting $\kappa$ greater than $\gamma$, meaning the short-term fluctuation factor converges faster than the long-term factor. The parameter $\mu_{\xi}\in\mathbb{R}$ represents the mean level of the long-term factor $\xi_{t}$, and we assume the short-term factor converges to zero. The parameters $\sigma_{\chi},\sigma_{\xi}\in\mathbb{R}^{+}$ are volatilities, representing the degree of variation in the price series over time. $\lambda_{\chi},\lambda_{\xi}\in\mathbb{R}$ are risk premiums. While we price commodities under the arbitrage-free assumption, in reality, mean term corrections, represented by $\lambda_{\chi}$ and $\lambda_{\xi}$, are necessary. $W_{t}^{\chi*}$ and $W_{t}^{\xi*}$ are correlated standard Brownian Motions with a correlation coefficient $\rho$. In discrete time, these processes are discretised to Gaussian noises, and all of these are parameters that need to be specified. Additionally, for simplicity, we assume that the standard errors $\sigma_{i},i=1,\dots,m$ for futures contracts are evenly spaced, i.e., $\sigma_{1}-\sigma_{2}=\sigma_{2}-\sigma_{3}=\dots=\sigma_{m-1}-\sigma_{m}$.
@@
Finally, all the simulated data can be downloaded. Click the “Download prices” and “Download maturities” buttons to download futures prices and maturities data, respectively. Note that although the Schwartz and Smith model specifies the logarithm of the spot price, all downloaded or plotted data are in real prices - they have been exponentiated. The “Generate new data” button allows users to simulate multiple realisations from the same set of parameters. Clicking on it will generate another set of random noises, resulting in different futures prices. This button is optional if users only need one realisation. The data will update automatically when you change any parameters.
@@ II. Implementation of the polynomial diffusion model
The procedure for simulating data from the polynomial diffusion model closely mirrors that of the Schwartz and Smith model, but it requires the specification of additional parameters.
@@
Both the polynomial diffusion model and the Schwartz and Smith model assume that the spot price $S_{t}$ is influenced by two latent factors, $\chi_{t}$ and $\xi_{t}$. However, while the Schwartz and Smith model assumes that the logarithm of $S_{t}$ is the sum of these two factors, the polynomial diffusion model posits that $S_{t}$ takes on a polynomial form. Currently, the PDSim GUI supports polynomials of degree 2, which can be expressed as:
@@
The coefficients $\alpha_{i},i=1,\dots,6$ are additional parameters specific to the polynomial diffusion model. If users wish to specify a polynomial of degree 1, they can simply set $\alpha_{4}=\alpha_{5}=\alpha_{6}=0$. Additionally, users are required to choose one of the non-linear filtering methods: the Extended Kalman Filter (EKF) or the Unscented Kalman Filter (UKF). All other procedures for simulating data follow the same steps as those in the Schwartz and Smith model.

### v2

@@ Abstract
PDSim is an R package that enables users to simulate commodity futures prices using the polynomial diffusion model introduced in Filipović and Larsson [11] through both a Shiny web application and R scripts. For user-supplied data, a standalone R routine has been developed to provide joint estimation of state variables and model parameters via the Extended Kalman Filter (EKF) or Unscented Kalman Filter (UKF). With its user-friendly interface, PDSim makes the features of simulations and estimations accessible. To date, it is the only package specifically designed for the simulation and estimation of the polynomial diffusion model. The Schwartz-Smith two-factor model [23] is also available within this package for both simulation and calibration.
The package is validated through several tests, including replication of the results in [23], unit testing of the coverage rate, and verification of the outputs of the main functions.
@@
Polynomial diffusion model; Schwartz-Smith model; commodity futures; two-factor commodity modelling; financial simulation
@@ Introduction
Stochastic models are crucial in the analysis of commodity futures, serving important roles in various financial areas such as price forecasting, risk management, and portfolio optimisation. Typically, the underlying spot price, denoted as $S_{t}$, is modelled as a function of certain factors. In 1990, the Ornstein-Uhlenbeck (OU) process was introduced to model oil futures in a two-factor setup, representing spot price and convenience yield [13]. Building on this foundation, Schwartz and Smith [23] modelled the logarithm of the underlying spot price of crude oil futures as the sum of two latent factors. These factors, assumed to follow the OU process, capture short-term fluctuations and the long-term equilibrium price level, respectively. This latent factor model and its extensions have been widely used in stochastic modelling in various fields, including biology [9], electricity forwards [18], agricultural commodity futures [24], and crude oil futures [2, 7, 22].
In this framework, the positivity of the futures price is ensured via an exponential transformation. However, in some markets, prices can become negative. For instance, in 2019, the natural gas price in the Permian Basin, West Texas, turned negative several times. On 20 April 2020, the front-month May 2020 WTI crude oil futures settled at –\$37.63 per barrel on the New York Mercantile Exchange, see, e.g., [10]. Negative prices occur even more frequently in the electricity market. In 2009, the province of Ontario, Canada, recorded 280 hours of negative electricity prices. In 2017, German electricity prices turned negative over 100 times. Between early 2018 and late 2020, electricity prices in Texas, USA, were negative for nearly 19% of total hours. Such cases illustrate that the previously mentioned framework is not applicable to these markets.
@@
The polynomial diffusion model relaxes this assumption. In this framework, the spot price is represented as a polynomial of any order in terms of the factors. In particular, under certain conditions, it can be proven that the conditional expectation of the spot price, which is equivalent to the futures price under the assumption of an arbitrage-free market and non-stochastic interest rate, is also a polynomial in terms of factors. The mathematical foundations of the polynomial diffusion model were introduced in [11], with applications of this framework seen in the modelling of electricity forwards, where the spot price is represented by a quadratic form of two factors [19].
In this paper, we introduce the R package PDSim, which is specifically designed to simulate futures prices and estimate latent factors using the polynomial diffusion model. PDSim consists of two components: the core package and a Shiny application of the same name. The core package provides exported functions for simulation as well as for linear and non-linear filtering. It is intended for users who wish to simulate data or fit real-world data. The Shiny app, by contrast, is designed exclusively for simulating data through a graphical user interface (GUI) and for visualising the results in a straightforward manner. This app has also been deployed on a Shiny server, allowing users to access it without requiring any local installation. This separation reflects the computational demands of parameter estimation when working with real-world data. Given the high dimensionality of the parameter space, estimation is computationally expensive, making R scripts a more suitable choice than a web-based interface in such cases.
In this paper, we provide a guide for using the GUI to simulate data and examples of how to use R scripts to both simulate and fit real-world data. The primary focus of PDSim is on simulation and joint estimation of state variables and model parameters, enabling users to explore the properties of the polynomial diffusion model. For applications involving real-world data, we recommend combining PDSim with established optimisation algorithms to minimise the negative log-likelihood returned by the KF, EKF, or UKF functions.
I. Schwartz-Smith two-factor model
@@
where $\chi_{t}$ represents the short-term fluctuation and $\xi_{t}$ is the long-term equilibrium price level. Additionally, we assume both $\chi_{t}$ and $\xi_{t}$ follow a risk-neutral Ornstein-Uhlenbeck (OU) process,
@@
In the original Schwartz-Smith model [23], $\gamma=0$, and only the short-term factor $\chi_{t}$ follows an OU process. Since 2000, many studies such as [2, 6, 20, 22] and others considered the extended version of the Schwartz-Smith model by introducing $\gamma\geq 0$. For simplicity, further, we continue to refer to this model with $\gamma\geq 0$ as the Schwartz-Smith model. Setting $\lambda_{\chi}=\lambda_{\xi}=0$ in Equation 2 and Equation 3 gives the “real processes” of the factors under statistical measure. The risk-neutral processes are used to calculate futures prices, and the real processes are for modelling state variables in real time.
@@
where
$x_{t}=\left[\begin{matrix}\chi_{t}\\ \xi_{t}\end{matrix}\right],c=\left[\begin{matrix}0\\ \frac{\mu_{\xi}}{\gamma}\left(1-e^{-\gamma\Delta t}\right)\end{matrix}\right],E=\left[\begin{matrix}e^{-\kappa\Delta t}&0\\ 0&e^{-\gamma\Delta t}\end{matrix}\right],$
$y_{t}=\left(\log{(F_{t,T_{1}})},\dots,\log{(F_{t,T_{m}})}\right)^{\top},d_{t}=\left(A(T_{1}-t),\dots,A(T_{m}-t)\right)^{\top},$
$F_{t}=\left[\begin{matrix}e^{-\kappa(T_{1}-t)},\dots,e^{-\kappa(T_{m}-t)}\\ e^{-\gamma(T_{1}-t)},\dots,e^{-\gamma(T_{m}-t)}\end{matrix}\right]^{\top}$
and $m$ is the number of futures contracts. The function $A(\cdot)$ is given by
@@
where $\tau$ is the time to maturity; $w_{t}$ and $v_{t}$ are multivariate Gaussian noise sequences with mean 0 and covariance matrix $\Sigma_{w}$ and $\Sigma_{v}$ respectively, where
@@ II. Polynomial diffusion model
Under the polynomial diffusion framework, the spot price $S_{t}$ is expressed as a polynomial function of the latent state vector $x_{t}$ (with components $\chi_{t}$ and $\xi_{t}$):
@@
In this context, we assume that the polynomial $p_{n}(x_{t})$ has a degree of 2 (or 1 if $\alpha_{4}=\alpha_{5}=\alpha_{6}=0$). However, it is worth noting that all the theorems presented here are applicable even for polynomials of higher degrees. Additionally, similar to the Schwartz-Smith model, we assume that $\chi_{t}$ and $\xi_{t}$ follow an Ornstein–Uhlenbeck process
@@
for $x\in\mathbb{R}^{d}$ and any $C^{2}$ function $f$. $\nabla^{2}$ and $\nabla$ denote the Hessian and Jacobian matrix of the function $f(\cdot)$, respectively, and $Tr(\cdot)$ denotes the trace of a matrix. Let $N$ be the dimension of $Pol_{n}$, and $H:\mathbb{R}^{d}\to\mathbb{R}^{N}$ be a function whose components form a basis of $Pol_{n}$. Then for any $p\in Pol_{n}$, there exists a unique vector $\vec{p}\in\mathbb{R}^{N}$ such that
@@
Obviously, the latent state vector $x_{t}$ satisfies the SDE with
@@
The Kalman Filter (KF) [14] is a commonly used filtering method in estimating latent state variables. However, KF can only deal with the linear Gaussian state model. To capture the non-linear dynamics in the polynomial diffusion model, we use Extended Kalman Filter (EKF) [16] and Unscented Kalman Filter (UKF) [17, 25]. Suppose we have a non-linear state-space model
@@ IV. Comparison with existing libraries
The graphical user interface (GUI) provides an easy way for anyone to use the PDSim package, even without programming knowledge. Simply enter the necessary parameters, and PDSim will simulate the data and generate well-designed interactive visualisations. Currently, PDSim supports data simulation from two models: the Schwartz-Smith two-factor model [23] and the polynomial diffusion model [11]. In this section, we first provide the guide to simulate data using GUI and R script. Then, we give an application of fitting WTI crude oil futures data using PDSim.
@@ I. Implementation of the Schwartz-Smith model
First, we establish some global configurations, as illustrated in Figure 3, including defining the number of observations (trading days) and contracts. Additionally, we select the model from which the simulated data will be generated.
Figure 4: Specify model parameters (Schwartz-Smith model).
For the Schwartz-Smith model, we assume that the logarithm of the spot price $S_{t}$ is the sum of two latent factors:
@@
Here, $\kappa,\gamma\in\mathbb{R}^{+}$ are the speed of mean-reversion parameters, controlling how quickly these two latent factors converge to their mean levels. Most experiments suggest that $\kappa$ and $\gamma$ fall within the range $(0,3]$. To avoid issues with parameter identification, we recommend setting $\kappa$ greater than $\gamma$, meaning the short-term fluctuation factor converges faster than the long-term factor. The parameter $\mu_{\xi}\in\mathbb{R}$ represents the mean level of the long-term factor $\xi_{t}$, and we assume the short-term factor converges to zero. The parameters $\sigma_{\chi},\sigma_{\xi}\in\mathbb{R}^{+}$ are volatilities, representing the degree of variation in the price series over time. The parameters $\lambda_{\chi},\lambda_{\xi}\in\mathbb{R}$ represent risk premiums. While we price commodities under the arbitrage-free assumption, in reality, mean term corrections, represented by $\lambda_{\chi}$ and $\lambda_{\xi}$, are necessary. $W_{t}^{\chi*}$ and $W_{t}^{\xi*}$ are correlated standard Brownian Motions with a correlation coefficient $\rho$. In discrete time, these processes are discretised to Gaussian noises, and all parameters are to be specified. Additionally, for simplicity, we assume that the standard errors $\sigma_{i},i=1,\dots,m$ for futures contracts are evenly spaced, i.e., $\sigma_{1}-\sigma_{2}=\sigma_{2}-\sigma_{3}=\dots=\sigma_{m-1}-\sigma_{m}$. All required parameters are shown in Figure 4.
@@
Finally, all the simulated data can be downloaded. Click the “Download prices” and “Download maturities” buttons to download the data for the futures prices and maturities, respectively. Note that although the Schwartz-Smith model specifies the logarithm of the spot price, all downloaded or plotted data are in real prices - they have been exponentiated. The “Generate new data” button allows users to simulate multiple realisations from the same set of parameters. Clicking on it will generate another set of random noises, resulting in different futures prices. This button is optional if users only need one realisation.Any change in parameters triggers an automatic data update. These three buttons are shown in Figure 5.
@@ II. Implementation of the polynomial diffusion model
The procedure for data simulation from the polynomial diffusion model closely mirrors that of the Schwartz-Smith model, but it requires the specification of additional parameters.
@@
Both the polynomial diffusion model and the Schwartz-Smith model assume that the spot price $S_{t}$ is the function of two latent factors, $\chi_{t}$ and $\xi_{t}$. However, while the Schwartz-Smith model assumes that the logarithm of $S_{t}$ is the sum of these two factors, the polynomial diffusion model assumes that $S_{t}$ takes on a polynomial form. Currently, the PDSim GUI supports polynomials of degree 2, which can be expressed as:
@@
The coefficients $\alpha_{i},i=1,\dots,6$ are additional parameters specific to the polynomial diffusion model. If users wish to specify a polynomial of degree 1, they can simply set $\alpha_{4}=\alpha_{5}=\alpha_{6}=0$. Users are also required to choose one of the non-linear filtering methods: the Extended Kalman Filter (EKF) or the Unscented Kalman Filter (UKF). All other procedures for data simulation follow the same steps as those in the Schwartz-Smith model. Additional parameters of the polynomial diffusion model are shown in Figure 6.
III. Implementation of the models using R script
The Shiny application simplifies the simulation and estimation procedures. However, if users require greater control over the simulated data, the R script can be used. This section demonstrates how to use scripts to simulate data through the polynomial diffusion model and estimate the latent state variables.
First, load PDSim and specify the global settings:
n_obs <- 100 # number of observations
n_contract <- 10 # number of contracts
dt <- 1/360 # interval between two consecutive time points,
# where 1/360 represents daily data
Next, specify the parameters and model coefficients:
# set of parameters
par <- c(0.5, 0.3, 1, 1.5, 1.3, -0.3, 0.5, 0.3,
seq(from = 0.1, to = 0.01, length.out = n_contract))
x0 <- c(0, 1/0.3) # initial values of state variables
n_coe <- 6 # number of model coefficient
par_coe <- c(1, 1, 1, 1, 1, 1) # model coefficients
Currently, PDSim supports a polynomial of order 2, which corresponds to six model coefficients. The next step is to define the measurement and state equations. We recommend using “state_linear” and “measurement_polynomial” within the package for a linear state equation and a polynomial measurement equation, respectively. However, users may also define their own functions:
# state equation
func_f <- function(xt, par) state_linear(xt, par, dt)
# measurement equation
func_g <- function(xt, par, mats)
measurement_polynomial(xt, par, mats, 2, n_coe)
Finally, simulate the data:
dat <- simulate_data(c(par, par_coe), x0, n_obs, n_contract,
func_f, func_g, n_coe, Gaussian”, 1234)
price <- dat\$yt # simulated futures price
mats <- dat\$mats # time to maturity
xt <- dat\$xt # state variables
The latent state variables can then be estimated using either EKF or UKF:
est_EKF <- EKF(c(par, par_coe, x0), price, mats,
func_f, func_g, dt, n_coe, Gaussian”)
est_UKF <- UKF(c(par, par_coe, x0), price, mats,
func_f, func_g, dt, n_coe, Gaussian”)
IV. Application to WTI crude oil futures data
We now demonstrate the implementation of PDSim on real-world data. Specifically, we use weekly WTI crude oil futures prices from 2 January 1990 to 14 February 1995, which is the exact same dataset used in [23]. This dataset is available in the R package “NFCP”.
For this implementation, we use the following code:
library(PDSim)
library(NFCP)
data(SS_oil)
contracts <- SS_oil\$stitched_futures
n_obs <- dim(contracts)[1]
n_contract <- dim(contracts)[2]
maturities <- matrix(rep(SS_oil\$stitched_TTM, n_obs),
nrow = n_obs, byrow = TRUE)
dt <- SS_oil\$dt
##### Schwartz Smith model #####
# state equation
func_f <- function(xt, par) state_linear(xt, par, dt)
# measurement equation
func_g <- function(xt, par, mats) measurement_linear(xt, par, mats)
# initial guesses
par0 <- c(1.5, 1.3, 2, 1.5, 1.5, 0.5, 1, 1, rep(0.1, n_contract))
x0 <- c(0, 0)
# negative log-likelihood
nll <- function(par) KF(par = par, yt = as.matrix(log(contracts)),
mats = maturities, delivery_time = 0,
dt = dt, smoothing = FALSE,
seasonality = None”)\$nll
# bounds
lb <- c(1e-04, 1e-04, -5, 1e-04, 1e-04, -0.9999, -5, -5,
rep(1e-04, n_contract), -5, -5) # lower bounds
ub <- c( 3, 3, 5, 3, 3, 0.9999, 5, 5,
rep( 1, n_contract), 5, 5) # upper bounds
# parameter estimation
result_SS <- optim(par = c(par0, x0),
fn = nll, method = L-BFGS-B”,
lower = lb, upper = ub)
# estimated futures contracts
est_SS <- KF(par = result_SS\$par, yt = as.matrix(log(contracts)),
mats = maturities, delivery_time = 0, dt = dt,
smoothing = FALSE, seasonality = None”)
yt_hat_SS <- data.frame(exp(func_g(t(est_SS\$xt_filter),
result_SS\$par,
maturities)\$y))
# mean absolute error of SS model
mae_SS <- colMeans(abs(log(contracts) - log(yt_hat_SS)))
##### Polynomial diffusion model #####
n_coe <- 6
# state equation
func_f <- function(xt, par) state_linear(xt, par, dt)
# measurement equation
func_g <- function(xt, par, mats)
measurement_polynomial(xt, par, mats, 2, n_coe)
# initial guesses
par0 <- c(1.5, 0.3, 0, 1.5, 1.5, 0, 0.5, 0.5, rep(0.1, n_contract))
par_coe0 <- c(1, 1, 1, 1, 1, 1)
x0 <- c(0, 0)
# negative log-likelihood
nll_EKF <- function(par) EKF(par = par, yt = as.matrix(contracts),
mats = maturities, func_f = func_f,
func_g = func_g, dt = dt, n_coe = n_coe,
noise = Gaussian”)\$nll
# bounds
lb <- c(1e-04, 1e-04, -5, 1e-04, 1e-04, -0.9999, -5, -5,
rep(1e-04, n_contract), rep(-5, n_coe), -5, -5) # lower bounds
ub <- c( 3, 3, 5, 3, 3, 0.9999, 5, 5,
rep( 1, n_contract), rep( 5, n_coe), 5, 5) # upper bounds
# parameter estimation
result_PD_EKF <- optim(par = c(par0, par_coe0, x0),
fn = nll_EKF, method = L-BFGS-B”,
lower = lb, upper = ub,
control = list(maxit = 10000))
# estimated futures contracts
est_EKF <- EKF(result_PD_EKF\$par, as.matrix(contracts),
maturities, func_f, func_g, dt, n_coe, Gaussian”)
yt_hat_EKF <- data.frame(func_g(t(est_EKF\$xt_filter),
result_PD_EKF\$par, maturities)\$y)
# mean absolute error of PD model
mae_EKF <- colMeans(abs(log(contracts) - log(yt_hat_EKF)))
Table 1 reports the mean absolute errors (MAEs) of the logarithm of futures prices estimated using both the Schwartz–Smith (SS) model and the polynomial diffusion (PD) model. For comparison, we also include the results from Schwartz and Smith’s original paper [23]. The MAEs obtained using PDSim for the SS model are comparable to those reported in the original study.
Table 1: Estimation mean absolute errors of the logarithm of futures prices with different maturities.
Maturity SS paper PDSim - SS model PDSim - PD model
1 month 0.0314 0.0268 0.0519
5 months 0.0035 0.0005 0.0279
9 months 0.0020 0.0030 0.0290
13 months 0.0000 0.0000 0.0338
17 months 0.0028 0.0038 0.0410
The performance of the polynomial diffusion model varies across markets and time periods. In the earlier example (Table 1), the Schwartz–Smith model provides more accurate estimates than the polynomial diffusion model for all contracts. However, Table 2 presents another application of PDSim to WTI crude oil futures data for the period 2015–2018. In this case, the polynomial diffusion model outperforms the Schwartz–Smith model, achieving a mean RMSE of 0.1921.
Table 2: Root mean square errors (RMSE) for each futures contract using the polynomial diffusion (PD) model and the Schwartz-Smith (SS) model for the period from 2015 to 2018.
Contracts PDSim - PD model PDSim - SS model
Contract 1 1.1481 0.8884
Contract 2 0.7283 0.5043
Contract 3 0.4262 0.2460
Contract 4 0.2522 0.1189
Contract 5 0.1347 0.0614
Contract 6 0.0516 0.0903
Contract 7 0.0136 0.1056
Contract 8 0.0323 0.1080
Contract 9 0.0385 0.0963
Contract 10 0.0251 0.0708
Contract 11 0.0018 0.0526
Contract 12 0.0322 0.0611
Contract 13 0.0693 0.0933
Mean 0.2272 0.1921
It is worth noting that fitting real data can be computationally expensive. This is primarily due to the high dimensionality of the parameter space, which leads to slow optimisation. For reference, fitting the WTI crude oil futures data using the polynomial diffusion model takes approximately 30 minutes. In this case, we employ the “L-BFGS-B” optimisation method introduced in [5]. We recommend that users experiment with different optimisation algorithms depending on the dataset, in order to achieve more efficient estimation.